OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)?

OpenStudy (anonymous):

A \[\frac{ - 3}{ 2 }\] B \[\frac{ - 2}{ 3}\] C \[\frac{ 3 }{ 2 }\] D \[\frac{ 2 }{ 3 }\]

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2) so when given points (-2, 4 ) and ( 2, -2) you get \[slope=\frac{ -2-4 }{ 2-(-2) } = ?\]

OpenStudy (anonymous):

\[\frac{ -3}{ 2 }\] ? ^u^

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Thank chu :3

jigglypuff314 (jigglypuff314):

you're welcome ^_^